# Dataset Distillation-based Hybrid Federated Learning on Non-IID Data

### v2

@@ Abstract
With the development of edge computing, Federated Learning (FL) has emerged as a promising solution for the intelligent Internet of Things (IoT). However, applying FL in mobile edge-cloud networks is greatly challenged by statistical heterogeneity and high communication overhead. To address it, we propose a hybrid federated learning framework called HFLDD, which integrates dataset distillation to generate approximately independent and equally distributed (IID) data, thereby improving the performance of model training. In particular, we partition the clients into heterogeneous clusters, where the data labels among different clients within a cluster are unbalanced while the data labels among different clusters are balanced. The cluster heads collect distilled data from the corresponding cluster members, and conduct model training in collaboration with the server. This training process is like traditional federated learning on IID data, and hence effectively alleviates the impact of non-IID data on model training. We perform a comprehensive analysis of the convergence behavior, communication overhead, and computational complexity of the proposed HFLDD. Extensive experimental results based on multiple public datasets demonstrate that when data labels are severely imbalanced, the proposed HFLDD outperforms the baseline methods in terms of both test accuracy and communication cost.
Index Terms:
Dataset distillation, hybrid federated learning, non-IID data, heterogeneous clusters
@@ I Introduction
In recent years, the rapid advancement of IoT devices and smart terminals has led to an exponential increase in global data volume [1]. Mobile edge computing (MEC) provides an efficient solution by offloading data storage, analysis, and processing to the network edge, significantly improving efficiency and response times [2]. Meanwhile, advances in deep neural networks have advanced tasks such as image classification [3, 4], object detection [5, 6], and semantic segmentation [7], making machine learning (ML) workloads a primary focus in MEC. To address challenges such as privacy leakage and high communication costs, Federated Learning (FL) has become an emerging machine learning paradigm and has attracted widespread attention [8, 9, 10].
@@
In the widely used FL-based method FedAvg [11], the server updates the global model by aggregating the local model parameters with simply weighted averaging. This method performs well when dealing with independently and identically distributed (IID) data. However, researchers have found that the data among clients is often non-IID in practice. In such cases, if the server simply aggregates the parameters of the local model with weighted averaging, the drift of the local model may cause the global model to deviate from the global optimum [12], leading to slower convergence of the global model.
@@
To mitigate the impact of non-IID data on FL-based model training, researchers have explored various approaches. At the algorithm level, efforts focus on improving loss functions [13, 14, 15] and aggregation methods [16, 17, 18]. However, they cannot fully alleviate the impact of non-IID data on model training and may have poor performance when dealing with high statistical heterogeneity. At the system level, client clustering techniques are introduced to create a multicenter framework [19, 20, 21]. These techniques aim to group clients with similar data distributions together to form clusters. However, this approach primarily focuses on the similarity of the client and does not fully exploit the potential advantages of the heterogeneity of the client. At the data level, data sharing is considered an effective strategy [22, 23, 24]. It mitigates the negative impact of non-IID data by sharing a small amount of local data. Although this method can improve the performance of the global model, the server typically lacks comprehensive knowledge of the client data distributions, making it difficult to obtain a uniformly distributed dataset. Moreover, this approach conflicts with the privacy protection principles of FL.
@@
Recent work on data collection shows promising potential to overcome restrictions on data sharing while protecting privacy. Dataset distillation[25, 26, 27, 28] aims to achieve good generalization performance by training models on a synthesized dataset much smaller than the real dataset. Ref. [29] indicates that the privacy of the transmission of distilled data should be no worse than the transmission of model parameters. Additionally, distilled data not only provides visual privacy, but also enhances the model’s resistance to membership inference attacks [30].
@@
Motivated by the above observations, we propose a new learning framework named HFLDD, integrating dataset distillation with hybrid federated learning (HFL). The key challenge of our research is how to effectively transmit distilled data among clients to compensate data heterogeneity. Specifically, the main idea of the design is to partition clients into a few clusters where data labels are heterogeneous within each cluster but balanced across different clusters. Furthermore, within each cluster, the cluster members transmit their distilled data to a cluster head. In addition, the adoption of client clustering and dataset distillation in FL can reduce the total communication cost of model training, while enhancing the generalization capabilities and performance of the model in diverse data environments.
@@
    Experimental results on public datasets demonstrate that the proposed HFLDD outperforms the baseline methods in terms of both test accuracy and communication cost, especially when the data labels of clients are highly unbalanced.
The remainder of this paper is organized as follows. Section II introduces the related work. Section III provides the preliminaries and problem statement. Section IV presents the proposed HFLDD framework in detail. Section V provides theoretical analysis on the convergence, communication cost, and computational complexity of HFLDD. Section VI conducts extensive experimental evaluations to validate the effectiveness of HFLDD. Section VII provides a discussion. Finally, Section VIII presents the conclusion.
@@ II RELATED WORK
II-A Federated Learning on non-IID data
The basic frameworks of FL include traditional FL, decentralized FL (DFL), and HFL. Addressing the performance decline of FL on non-IID data is a critical challenge. Each framework employs different methods to address this problem.
In traditional FL, all clients train models locally and then transmit these local models to a central server for global aggregation. Various methods have been proposed since the introduction of FedAvg [11] to address this heterogeneity issue. One main direction attempts to improve the design of the local loss function. FedProx [14] introduces an additional $L_{2}$ regularization term in the loss function to limit the size of local updates. SCAFFOLD [13] applies the variance reduction technique to correct local training. MOON [15] uses the similarity between model representations to correct local client training. However, when the data heterogeneity is severe, the accuracy achieved remains low for regularization-based and weight-adjustment methods.
@@ II-A Federated Learning on non-IID data
DFL can implement distributed learning from data distributed across multiple clients without the need for a central server. To address data heterogeneity issues in DFL, existing research focused primarily on how to utilize the topological structure to improve the learning process. D-Cliques[31] address label shift problems in decentralized settings by constructing heterogeneous topological structures based on the underlying data distribution of each client, which will induce privacy leakage. In consideration of privacy preservation, Ref. [32] applies a proxy-based method to learn the knowledge of local dataset labels based on a global dataset and to construct a heterogeneous topological structure for DFL. The proposed method has shown that when the neighbors of each client have heterogeneous data, it can improve the convergence of model training in addressing label distribution skew problems. This insight inspires the main idea of our method. It should be noted that the paradigm of fully distributed topological structure construction and learning suffers from a high communication overhead.
@@
Unlike traditional and decentralized FL, HFL involves both client-to-server communication and client-to-client communication. In [33], a multistage HFL algorithm is developed to handle the issue of non-IID data by continuously adjusting the rounds of client-to-client communication and global aggregation cycles. In particular, Ref. [34] proposes an HFL framework, called FedSeq, based on client clustering and sequential training. It partitions clients into multiple clusters, with each cluster head communicating with the server. Within each cluster, clients form a ring topology for sequential training. The core idea of FedSeq is that the model in the cluster head is based on the training data of all the clients within this cluster, thereby improving the test accuracy. FedSeq requires multiple rounds of parameter transmission among clients, which still incurs significant communication cost.
II-B Dataset Distillation
The current dataset distillation frameworks are mainly divided into meta-learning frameworks and data matching frameworks [35]. Ref. [25] proposes a meta-learning framework that adopts a dual optimization approach. However, this method involves a nested loop that can be difficult to optimize. Considering the existence of closed-form solutions in the kernel regression paradigm, Ref. [36] replaces the neural network in the inner loop of the meta-learning framework with a kernel model. In the data matching framework, DC [37, 38] and DM [39, 40] aim to match the gradients and distribution of the distilled data with the original data during training. Trajectory matching [41] generates synthetic data by matching training trajectories.
Recently, the integration of dataset distillation into FL has attracted widespread attention. In DynaFed[42], the server generates distilled data by matching the trajectory after each aggregation and further updates the global model based on these data. Ref. [43] improves communication efficiency and test accuracy by constructing distilled data on each client to locally match the loss landscape of the original data. However, these works do not consider the use of D2D (device-to-device) networks to compensate for client heterogeneity. Ref. [44] introduces a method for extracting meta-knowledge from clients and exchanging it among clients, but this knowledge sharing may be less effective in scenarios with severe data heterogeneity.
In comparison with the aforementioned works, we design a new framework of HFL to address the issue of non-IID data. It extends the idea of constructing heterogeneous topological structure in DFL to HFL, and fully leverages the potential of heterogeneity among clients.
@@ III-A Dataset Distillation
The main idea of dataset distillation is to transfer a large dataset to a small dataset, while preserving the essential characteristics of the original dataset. Suppose that the original data set and the distilled dataset are, respectively, denoted by $\mathcal{T}=\left\{\left(\mathrm{x}_{t,i},y_{t,i}\right)\right\}_{i=1}^{|\mathcal{T}|}$ and $\mathcal{S}=\left\{\left(\mathrm{x}_{s,j},y_{s,j}\right)\right\}_{j=1}^{|\mathcal{S}|}$, where $\mathrm{x}_{t,i}\in\mathbb{R}^{d}$ and $y_{t,i}\in\mathbb{R}^{n_{c}}$ are, respectively, the input data and the corresponding label of the $i$-th sample in the original dataset. $\mathrm{x}_{s,j}\in\mathbb{R}^{d}$ and $y_{s,j}\in\mathbb{R}^{n_{c}}$ are, respectively, the input data and the corresponding label of the $j$-th sample in the distilled dataset. $|\mathcal{T}|$ and $|\mathcal{S}|$ represent the sizes of the original dataset and the distilled dataset, respectively. Generally, the size of $\mathcal{T}$ is considerably larger than $\mathcal{S}$, i.e., $|\mathcal{S}|\ll|\mathcal{T}|$.
To realize dataset distillation, one of the most widely used methods is Kernel Inducing Points (KIP) [45, 36], which is a first-order meta-learning algorithm based on kernel ridge regression (KRR). Specifically, the KIP method is to find a synthetic dataset $\mathcal{S}$ that is approximate to the target dataset $\mathcal{T}$ by minimizing the following loss function
$\mathcal{L}(\mathcal{S})=\frac{1}{2}\left\|y_{t}-K_{X_{t}X_{s}}\left(K_{X_{s}X_{s}}+\lambda I\right)^{-1}y_{s}\right\|^{2},$ (1)
where $\lambda>0$ is a regularization parameter. $y_{t}\in\mathbb{R}^{|\mathcal{T}|\times n_{c}}$ and $y_{s}\in\mathbb{R}^{|\mathcal{S}|\times n_{c}}$ are the labels of $\mathcal{T}$ and $\mathcal{S}$, respectively. ${K}_{X_{s}X_{s}}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$ and $K_{X_{t}X_{s}}\in\mathbb{R}^{|\mathcal{T}|\times|\mathcal{S}|}$ are two kernel matrices defined as $(K_{X_{s}X_{s}})_{ij}=k\left(\mathrm{x}_{s,i},\mathrm{x}_{s,j}\right)$ and $(K_{X_{t}X_{s}})_{ij}=k\left(\mathrm{x}_{t,i},\mathrm{x}_{s,j}\right)$, where $k(\cdot)$ denotes the kernel function. KIP is highly efficient for the distillation of datasets because it involves only first-order optimization.

### v3

@@ Abstract
In federated learning, the heterogeneity of client data has a great impact on the performance of model training. Many heterogeneity issues in this process are raised by non-independently and identically distributed (non-IID) data. To address the issue of label distribution skew, we propose a hybrid federated learning framework called HFLDD, which integrates dataset distillation to generate approximately independent and equally distributed (IID) data, thereby improving the performance of model training. In particular, we partition the clients into heterogeneous clusters, where the data labels among different clients within a cluster are unbalanced while the data labels among different clusters are balanced. The cluster heads collect distilled data from the corresponding cluster members, and conduct model training in collaboration with the server. This training process is like traditional federated learning on IID data, and hence effectively alleviates the impact of non-IID data on model training. We perform a comprehensive analysis of the convergence behavior, communication overhead, and computational complexity of the proposed HFLDD. Extensive experimental results based on multiple public datasets demonstrate that when data labels are severely imbalanced, the proposed HFLDD outperforms the baseline methods in terms of both test accuracy and communication cost.
@@ I Introduction
In recent years, the rapid advancement of IoT devices and smart terminals has led to an exponential increase in global data volume [yuan2024secure]. Mobile edge computing (MEC) provides an efficient solution by offloading data storage, analysis, and processing to the network edge, significantly improving efficiency and response times [li2018learning, zhao2022energy]. Meanwhile, advances in deep neural networks have advanced tasks such as image classification [krizhevsky2012imagenet, han2022survey], object detection [redmon2018yolov3, ren2016faster, wang2024end], and semantic segmentation [long2015fully], making machine learning (ML) workloads a primary focus in MEC. To address challenges such as privacy leakage and high communication costs, Federated Learning (FL) has become an emerging machine learning paradigm and has attracted widespread attention [cai2025comprehensive, han2025rethinking, yin2021comprehensive, li2025accelerating].
@@
In the widely used FL-based method FedAvg [mcmahan2017communication], the server updates the global model by aggregating the local model parameters with simply weighted averaging. This method performs well when dealing with IID data. However, researchers have found that the data among clients is often non-IID in practice[kairouz2021advances]. In such cases, if the server simply aggregates the parameters of the local model with weighted averaging, the drift of the local model may cause the global model to deviate from the global optimum [li2019convergence], leading to slower convergence of the global model[wang2021field].
@@
To mitigate the impact of non-IID data on FL-based model training, researchers have explored various approaches. At the algorithm level, efforts focus on improving loss functions [karimireddy2020scaffold, li2020federated, li2021model, guo2025FedVLS, lu2023FedLMD, lee2022FedNTD, acar2021FedDyn, zhang2022FedLC] and aggregation methods [tan2022adafed, chai2021fedat, ye2023feddisco, li2023fedlaw, rehman2023Ldawa, shi2025fedawa]. However, they cannot fully alleviate the impact of non-IID data on model training and may have poor performance when dealing with high statistical heterogeneity. At the system level, client clustering techniques are introduced to create a multicenter framework [ghosh2020efficient, he2022improving, wang2021edge]. These techniques aim to group clients with similar data distributions together to form clusters. However, this approach primarily focuses on the similarity of the client and does not fully exploit the potential advantages of the heterogeneity of the client. At the data level, data sharing is considered an effective strategy [zhao2018federated, gu2022fedaux, tuor2021overcoming]. It mitigates the negative impact of non-IID data by sharing a small amount of local data. Although this method can improve the performance of the global model, the server typically lacks comprehensive knowledge of the client data distributions, making it difficult to obtain a uniformly distributed dataset. Moreover, this approach conflicts with the privacy protection principles of FL.
@@
Recent work on dataset distillation shows promising potential to overcome restrictions on data sharing while protecting privacy [10273632, song2023federated, liu2022meta, pi2023dynafed, jia2024unlocking, xiong2023feddm, xu2024flip, wang2024aggregation, yan2025fedvck]. Dataset distillation[wang2018dataset, nguyen2021dataset, zhao2020dataset, zhao2023dataset, cazenavette2022dataset] aims to achieve good generalization performance by training models on a synthesized dataset that is much smaller than the real dataset. It has been indicated that the privacy of the transmission of distilled data should not be worse than that of the transmission of model parameters[zhou2020distilled]. Additionally, distilled data not only provides visual privacy, but also enhances the model’s resistance to membership inference attacks [dong2022privacy].
@@
Motivated by the above observations, we propose a new learning framework named HFLDD, integrating dataset distillation with hybrid federated learning (HFL). The main idea of the design is to partition clients into a few clusters where data labels are heterogeneous within each cluster but balanced across different clusters. Furthermore, within each cluster, the cluster members transmit their distilled data to a cluster head. The adoption of client clustering and dataset distillation in FL can reduce the total communication cost of model training, while compensating for data heterogeneity between clients and improving the performance of the global model.
@@
    Experimental results in public datasets demonstrate that the proposed HFLDD outperforms baseline methods in terms of both test accuracy and communication cost, especially when the client data labels are highly unbalanced.
The remainder of this paper is organized as follows. Section II introduces the related work. Section III provides the preliminaries and problem statement. Section IV presents the proposed HFLDD framework in detail. Section V provides a theoretical analysis on the convergence, communication cost, and computational complexity of HFLDD. Section VI conducts extensive experimental evaluations to validate the effectiveness of HFLDD. Section VII provides a discussion. Finally, Section VIII presents the conclusion.
@@ II RELATED WORK
Non-IID data poses a significant challenge to FL. Typical non-IID scenarios are generally classified into three categories [kairouz2021noniidscenary]: feature distribution skew, quantity skew, and label distribution skew. Among these, label distribution skew, can be further divided into quantity-based and distribution-based subtypes. To address the non-IID issues, researchers have explored solutions from the algorithm, system, and data levels. Specifically, algorithm level methods concentrate on improving loss functions and aggregation methods; system-level methods focus on optimizing the framework’s topological structure; and data-level methods seek to improve sharing, augmentation and distillation of local data. The basic frameworks of FL include traditional FL, decentralized FL (DFL), and HFL. Each framework employs different methods to address the non-IID problem.
II-A Traditional FL on Non-IID data
In traditional FL, all clients train models locally and then transmit these local models to a central server for global aggregation. Various methods have been proposed since the introduction of FedAvg [mcmahan2017communication] to address the non-IID issue. Approaches such as FedProx [li2020federated], SCAFFOLD [karimireddy2020scaffold], and MOON [li2021model] improve local training by modifying loss functions or introducing representation regularization. Another line of work focuses on dynamically adjusting aggregation weights. FedDisco [ye2023feddisco] determines weights by jointly considering dataset scale and discrepancies between local and global category distributions, while FedLaw [li2023fedlaw] adaptively refines aggregation weights through gradient descent on a proxy dataset at the server, enabling a more balanced and data-aware global aggregation.
Apart from these model-centric approaches, data-centric solutions have also been explored, including data augmentation [yoon2021fedmix], data selection [cao2023knowledge], and data sharing [zhao2018federated]. Recently, dataset distillation has emerged as a promising solution for data-centric federated learning. FedMK [liu2022meta] proposes dynamic weight assignment and meta-knowledge sharing to accelerate convergence. DynaFed [pi2023dynafed] generates distilled data by matching the global model trajectory after each aggregation and further updates the global model based on these distilled data. FedDM [xiong2023feddm] improves communication efficiency and test accuracy by constructing distilled data on each client to locally match the loss landscape of the original data. FLiP [xu2024flip] applied the principle of least privilege to ensure that clients only share the necessary knowledge, further protecting sensitive information. FedAF [wang2024aggregation] enables clients to collaboratively learn distilled data and soft labels to improve model performance. FedVCK [yan2025fedvck] condenses valuable client knowledge into distilled data with latent distribution constraints, improving non-IID robustness and communication efficiency.
II-B Decentralized FL on Non-IID data
@@
DFL can implement distributed learning from data distributed across multiple clients without the need for a central server. DESA [huang2024DESA] employs synthetic anchor data to achieve implicit feature alignment and cross-client knowledge transfer without relying on a central coordinator. DFPL [zhang2025dfpl] enables client collaboration through prototype exchange and integrates blockchain to support a decentralized architecture. Furthermore, some research focused primarily on how to utilize the topological structure to improve the learning process. DFedPGP [liu2024DFedPGP] leverages directed communication topologies and gradient push optimization to efficiently learn shared feature layers and personalized classifier heads. D-Cliques[bellet2022d] address label shift problems by constructing heterogeneous topological structures based on the underlying data distribution of each client, which may induce privacy leakage. To enhance privacy preservation, Ref. [abebe2023optimizing] adopts a proxy-based method to learn the knowledge of local dataset labels based on a global dataset and to construct a heterogeneous topological structure for DFL. The proposed method has shown that when the neighbors of each client have heterogeneous data, it can improve the convergence of model training in addressing label distribution skew problems. This insight inspires the main idea of our method. It should be noted that the paradigm of fully distributed topological structure construction and learning suffers from a high communication overhead and node inconsistency.
TABLE I: Comparisons of different FL frameworks on Non-IID data
Framework Reference Non-IID scenario Method Limitation
Algorithm
level
System
level
Data
level
Traditional FL FedAvg[mcmahan2017communication] Quantity-based label distribution skew Sensitive to data heterogeneity; Server coordination required.
FedProx[li2020federated] Quantity-based label distribution skew
SCAFFOLD[karimireddy2020scaffold] Distribution-based label distribution skew
MOON[li2021model] Distribution-based label distribution skew
FedDisco[ye2023feddisco] Label distribution skew
FedLaw[li2023fedlaw] Distribution-based label distribution skew
FedMK[liu2022meta] Distribution-based label distribution skew
DynaFed[pi2023dynafed] Distribution-based label distribution skew
FedDM[xiong2023feddm] Distribution-based label distribution skew
Decentralized FL DESA[huang2024DESA] Feature distribution skew High communication overhead; Node inconsistency.
DFPL[zhang2025dfpl] Quantity-based label distribution skew
DFedPGP[liu2024DFedPGP] Label distribution skew
D-Cliques[bellet2022d] Quantity-based label distribution skew
Ref.[abebe2023optimizing] Label distribution skew
Hybrid FL HFML[li2025hfml] Quantity-based label distribution skew High communication overhead.
Semi-FL[chen2020semifl] Quantity-based label distribution skew
FedCluster[chen2020fedcluster] Distribution-based label distribution skew
Ref.[lin2021semi] Quantity-based label distribution skew
FedSeq[chen2023fedseq] Distribution-based label distribution skew
II-C Hybrid FL on Non-IID data
@@
Unlike traditional and decentralized FL, HFL involves both client-to-server communication and client-to-client communication. Many existing approaches focus on system-level innovations by designing novel architectures. For example, HFML [li2025hfml] introduces a client–edge–cloud architecture that applies deep mutual learning at the client–edge layer for knowledge sharing. In [lin2021semi], a multistage HFL algorithm is developed to handle non-IID data by continuously adjusting the rounds of client-to-client communication and global aggregation cycles. Another representative system-level direction is client clustering. Semi-FL [chen2020semifl] clusters clients and adopts in-cluster sequential training, enabling neighboring clients to share models and reduce communication overhead. FedCluster [chen2020fedcluster] accelerates convergence by clustering clients and performing cyclic training. FedSeq [chen2023fedseq] is based on client clustering and sequential training. It partitions clients into multiple clusters, with each cluster head communicating with the server. Within each cluster, clients form a ring topology for sequential training. The core idea of FedSeq is that the model in the cluster head is based on the training data of all the clients within this cluster, thereby improving the test accuracy.
Table I summarizes different FL frameworks on non-IID data. In comparison with the aforementioned works, we design a new HFL framework to mitigate the challenges of non-IID data from both system level and data level. By constructing heterogeneous topology among clients and incorporating dataset distillation, the proposed method can effectively compensate the data heterogeneity among clients and achieve a well-performed global model with low communication overhead.
@@ III-A Dataset Distillation
The main idea of dataset distillation is to transfer a large dataset to a small dataset, while preserving the essential characteristics of the original dataset. Suppose that the original data set and the distilled data are, respectively, denoted by $\mathcal{T}=\left\{\left({x}_{i},y_{i}\right)\right\}_{i=1}^{|\mathcal{T}|}$ and $\mathcal{S}=\left\{\left(\hat{{x}}_{j},\hat{y}_{j}\right)\right\}_{j=1}^{|\mathcal{S}|}$, where $\mathrm{x}_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}^{n_{c}}$ are, respectively, the input data and the corresponding label of the $i$-th sample in the original dataset; $\hat{{x}}_{j}\in\mathbb{R}^{d}$ and $\hat{y}_{j}\in\mathbb{R}^{n_{c}}$ are, respectively, the input data and the corresponding label of the $j$-th sample in the distilled data; $|\mathcal{T}|$ and $|\mathcal{S}|$ represent the sizes of the original dataset and the distilled data, respectively. Generally, the size of $\mathcal{T}$ is considerably larger than $\mathcal{S}$, i.e., $|\mathcal{S}|\ll|\mathcal{T}|$.
To realize dataset distillation, one of the most widely used methods is Kernel Inducing Points (KIP) [nguyen2020dataset, nguyen2021dataset], which is a first-order meta-learning algorithm based on kernel ridge regression (KRR). Specifically, the KIP method is to find a ditilled dataset $\mathcal{S}$ that is approximate to the target dataset $\mathcal{T}$ by minimizing the following loss function
$\mathcal{L}(\mathcal{S})=\frac{1}{2}\left\|Y_{\mathcal{T}}-K_{X_{t}X_{s}}\left(K_{X_{s}X_{s}}+\lambda I\right)^{-1}Y_{\mathcal{S}}\right\|^{2},$ (1)
where $\lambda>0$ is a regularization parameter; $Y_{\mathcal{T}}\in\mathbb{R}^{|\mathcal{T}|\times n_{c}}$ and $Y_{\mathcal{S}}\in\mathbb{R}^{|\mathcal{S}|\times n_{c}}$ are the labels of $\mathcal{T}$ and $\mathcal{S}$, respectively; ${K}_{X_{s}X_{s}}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$ and $K_{X_{t}X_{s}}\in\mathbb{R}^{|\mathcal{T}|\times|\mathcal{S}|}$ are two kernel matrices defined as $(K_{X_{s}X_{s}})_{ij}=k\left(\hat{x}_{i},\hat{x}_{j}\right)$ and $(K_{X_{t}X_{s}})_{ij}=k\left({x}_{i},\hat{x}_{j}\right)$, where $k(\cdot)$ denotes the kernel function. KIP is highly efficient for the distillation of datasets because it involves only first-order optimization.